OpenStudy (anonymous):

solve by factoring 69p^2-1)=5p

OpenStudy (cj49):

69p^2-5p-1 use the same formulea

OpenStudy (anonymous):

isnt the formula ax^2+bx+c=0

OpenStudy (cj49):

yes instead of x its p

OpenStudy (anonymous):

so the first step is 6(p^2-1=5p 6(p^2= 5p+1

OpenStudy (cj49):

no it wud be 69p^2-5p-1 a=69 b=-5 c=-1

OpenStudy (anonymous):

its supposed to be 6 and then a parenthesis not a 9 sorry lol

OpenStudy (cj49):

6p^2-5p-6

OpenStudy (anonymous):

where did you get the -6 from?

OpenStudy (cj49):

6(p^2-1)=6p^2-6

OpenStudy (anonymous):

what do I do next?

OpenStudy (cj49):

a=6,b=-5 and c=-6

OpenStudy (anonymous):

whats the next step?

OpenStudy (cj49):

plug in the values in the formulae

OpenStudy (cj49):

p=3/2 , -2/3

OpenStudy (anonymous):

got it(:

OpenStudy (cj49):

cool....

OpenStudy (anonymous):

my college algebra class makes my head hurt lol